OpenStudy (anonymous):

Which equations are represented by the graph? Choose all answers that are correct.

OpenStudy (anonymous):

On the graph the line is on (2,-4) 4x – 2y = 8 y – 4 = 2x y = 2x – 4 y = 1/2x - 4